OpenStudy (anonymous):

A bag contains 6 red marbles, 4 blue marbles, 7 green marbles, and 3 yellow marbles. A marble is drawn from the bag and is not replaced. Then a second marble is drawn. What is the probability that both marbles drawn were red?

OpenStudy (anonymous):

.

OpenStudy (anonymous):

the probability that the first one is red, times the probability that the second one is red given that the first one is red

OpenStudy (anonymous):

45/637 then? Thank you!

OpenStudy (anonymous):

i didn't check, but it should be \[\frac{6}{20}\times \frac{5}{19}\]

OpenStudy (anonymous):

which is definitely not \(\frac{45}{637}\)

OpenStudy (anonymous):

ok it's .0789473684

OpenStudy (anonymous):

which is 3/38?

OpenStudy (anonymous):

\[\frac{6}{20}\times \frac{5}{19}=\frac{6}{4\times 19}=\frac{3}{38}\]

OpenStudy (anonymous):

yeah that is it

OpenStudy (anonymous):

ok, thank you so much!